ec 16 STATISTICS BOOK MAT 2 Micro 31. If the supply curve is given by the equation P- 10+ 40, and the demand curve is given by the equation P = 185-100, then a price ceiling set at Pmax-$50will result in a dead-weight loss of D:59:rr.oo 3:P-10叫(rs.り P.pdQ A) $45.00 B) $20.50 C) $43.75 D) $40.25 The equation for the demand curve is P--685-(2)Q. When Q goes from| 36 to 137, then the price must go from A) 413; 411 B) 414; 413 C) 413;414 D) 414; 411 32· to 33. If the supply curve is given by the equation P- 10+ 40, and the competitive market equilibrium price is $60, then a price-ceiling set at Pmax $50 will result in a producer surplus o A) $400 B) $300 C) $500 D) $200
Answer #1

31. P=10+4Q

P= 185-10Q

At equilibrium , demand and supply are equal.

10+4 Q= 185-10Q

14Q= 175

Q= 12.5 (Equilibrium quantity)

P= 10+(4)(12.5)

P= 60 (Equilibrium price)

At Price ceiling of P=50

P= 10+4Q

50= 10+4Q

Q= 10 (quantity supplied after price ceiling)

P=185-10Q

50= 185-10Q

Q= 13.5 (quantity demanded after price ceiling)

When Q=10, then P=185-10(10)= 85.

Deadweight loss = (0.5)(85-60)(12.5-10) + (0.5)(60-50)(12.5-10)

=(0.5)(25)(2.5) +(0.5)(10)(2.5)

= 31.25 +12.5 =$ 43.75

Hence, option(C) is correct.

32. Demand equation:P=685-2Q

When Q=136 , P=685-2(136)= $413

When Q=137 , P=685-2(137)= $ 411

When Q goes from 136 to 137 ,then P must goes from 413 to 411. Hence, option(A) is correct.

33. P=10+4Q

When Q=0 , P=10

When P=50 , 50=10+4Q

Q= 10

So, the producer surplus = (0.5)(50-10)(10)= $ 200

Hence, option(D) is correct.